[The following information applies to the questions displayed below.]

The following unadjusted trial balance is prepared at fiscal year-end for Nelson Company.

NELSON COMPANY
Unadjusted Trial Balance
January 31, 2017
Debit Credit
Cash $ 35,750
Merchandise inventory 14,500
Store supplies 5,300
Prepaid insurance 2,800
Store equipment 42,600
Accumulated depreciation—Store equipment $ 18,250
Accounts payable 17,000
Common stock 3,000
Retained earnings 15,000
Dividends 2,000
Sales 147,150
Sales discounts 1,850
Sales returns and allowances 2,100
Cost of goods sold 38,000
Depreciation expense—Store equipment 0
Salaries expense 31,600
Insurance expense 0
Rent expense 14,000
Store supplies expense 0
Advertising expense 9,900
Totals $ 200,400 $ 200,400

Rent expense and salaries expense are equally divided between selling activities and general and administrative activities. Nelson Company uses a perpetual inventory system.

Additional Information:

  1. Store supplies still available at fiscal year-end amount to $1,950.
  2. Expired insurance, an administrative expense, for the fiscal year is $1,750.
  3. Depreciation expense on store equipment, a selling expense, is $1,700 for the fiscal year.
  4. To estimate shrinkage, a physical count of ending merchandise inventory is taken. It shows $10,900 of inventory is still available at fiscal year-end.

Required:

1. Using the above information prepare adjusting journal entries:
2. Prepare a multiple-step income statement for fiscal year 2017.
3. Prepare a single-step income statement for fiscal year 2017.

Prepare a single-step income statement for fiscal year 2017.

NELSON COMPANY
Income Statement
For Year Ended January 31, 2017
Expenses
Total expenses 0
Answer #1

1. Preparing Adjusting Entries:-

Nelson Company

General Journal (Adjusting Entries)

For the Year Ended January 31,2017

Date Accounts Title Debit Credit
January 31 (a) Store Supplies Expenses $3,350
Store Supplies $3,350
January 31 (b) Insurance Expenses $1,750
Prepaid Insurance $1,750
January 31 (c) Depreciation Expenses $1,700
Accumulated Depreciation $1,700
January 31 (d) Cost of Goods Sold $3,600
Merchandise Inventory $3,600

Note:-

Store Supplies Expenses=(Store Supplies- Store Supplies at the end)

=$(5,300-1,950)=$3,350

Cost of Goods Sold=(Merchandise Inventory- Merchandise Inventory at the end)

=$(14,500-10,900)=$3,600

Before Preparing Income Statement we need to prepare Adjusted Trial Balance:-

Nelson Company

Worksheet

As on January 31,2017

Accounts Title Unadjusted Trial Balance Adjustments Adjusted Trial Balance
Debit Credit Debit Credit Debit Credit
Cash $35,750 $35,750
Merchandise Inventory 14,500 3,600 10,900
Store Supplies 5,300 3,350 1,950
Prepaid Insurance 2,800 1,750 1,050
Store Equipment 42,600 42,600
Accumulated Depreciation- Store Equipment $18,250 1,700 19,950
Accounts Payable 17,000 17,000
Common Stock 3,000 3,000
Retained Earnings 15,000 15,000
Dividends 2,000 2,000
Sales 147,150 147,150
Sales Discounts 1,850 1,850
Sales Returns and Allowances 2,100 2,100
Cost of Goods Sold 38,000 3,600 41,600
Depreciation Expenses- Store Equipment 0 1,700 1,700
Salaries Expenses 31,600 31,600
Insurance Expenses 0 1,750 1,750
Rent Expenses 14,000 14,000
Store Supplies Expenses 0 3,350 3,350
Advertising Expenses 9,900 9,900
Totals $200,400 $200,400 $10,400 $10,400 $202,100 $202,100

2. Preparing Multi Step Income Statement:-

Nelson Company

Multi Step Income Statement

For the Year Ended January 31,2017

Accounts Title Amount Amount
Revenue:-
Sales $147,150
Less:- Sales Discounts (1,850)
Less:- Sales Returns and Allowances (2,100)
Net Sales $143,200
Less:- Cost of Goods Sold (41,600)
Gross Profit $101,600
Expenses:-
Selling Expenses:-
Depreciation Expenses- Store Equipment 1,700
Salaries Expenses 15,800
Rent Expenses 7,000
Store Supplies Expenses 3,350
Advertising Expenses 9,900
Total Selling Expenses $37,750
Administrative Expenses:-
Rent Expenses 7,000
Salaries Expenses 15,800
Insurance Expenses 1,750
Total Administrative Expenses $24,550
Total Expenses $(62,300)
Net Income $39,300

Note:-

Rent and Salaries Expenses Equally divided between Selling and Administrative Expenses.

Rent Expenses for selling=$(14,000/2)=$7,000

Rent Expenses for Administrative Expenses=$(14,000/2)=$7,000

Salaries Expenses for Selling Expenses=$(31,600/2)=$15,800

Salaries Expenses for Administrative Expenses=$(31,600/2)=$15,800

3. Preparing Single Step Income Statement:-

Nelson Company

Single Step Income Statement

For the Year Ended January 31,2017

Accounts Title Amount Amount
Revenue:-
Sales $147,150
Less:- Sales Discounts (1,850)
Less:- Sales Returns and Allowances (2,100)
Net Sales $143,200
Less:- Cost of Goods Sold (41,600)
Gross Profit $101,600
Expenses:-
Depreciation Expense- Store Equipment 1,700
Salaries Expenses 31,600
Insurance Expenses 1,750
Rent Expenses 14,000
Store Supplies Expenses 3,350
Advertising Expenses 9,900
Total Expenses $(62,300)
Net Income $39,300